OpenStudy (fanduekisses):

Does 3x^2 times sqrt of x simply equals 3x^2sqrt(x) or...

OpenStudy (fanduekisses):

Should I treat the sqrt of x as x^1/2 ? If so would it equal 3x^2.5 ?

OpenStudy (gorv):

\[3x^2*\sqrt{2}\]

ganeshie8 (ganeshie8):

You're right! \[\large 3x^2 \sqrt{x} = 3x^2\cdot x^{\frac{1}{2}} = 3x^{2+\frac{1}{2}} = 3x^{\frac{5}{2}}\]

OpenStudy (gorv):

when base is same and number is multiplied than power will be added

OpenStudy (gorv):

here base is x so power of x will be added

OpenStudy (gorv):

2+1/2=2.5 so 3x^2.5

OpenStudy (fanduekisses):

ooohh great yay! It feels so good to understand something in math hehe thanks guys :)